# Partial Diacritization: A Context-Contrastive Inference Approach

### v2

@@ Abstract
Diacritization plays a pivotal role in improving readability and disambiguating the meaning of Arabic texts. Efforts have so far focused on marking every eligible character (Full Diacritization). Comparatively overlooked, Partial Diacritzation (PD) is the selection of a subset of characters to be marked to aid comprehension where needed. Research has indicated that excessive diacritic marks can hinder skilled readers—reducing reading speed and accuracy. We conduct a behavioral experiment and show that partially marked text is often easier to read than fully marked text, and sometimes easier than plain text. In this light, we introduce Context-Contrastive Partial Diacritization (CCPD)—a novel approach to PD which integrates seamlessly with existing Arabic diacritization systems. CCPD processes each word twice, once with context and once without, and diacritizes only the characters with disparities between the two inferences. Further, we introduce novel indicators for measuring partial diacritization quality {SR, PDER, HDER, ERE}, essential for establishing this as a machine learning task. Lastly, we introduce TD2, a Transformer-variant of an established model which offers a markedly different performance profile on our proposed indicators compared to all other known systems.111Demo: huggingface.co/spaces/bkhmsi/Partial-Arabic-Diacritization
Partial Diacritization:
A Context-Contrastive Inference Approach
@@ 1 Introduction
There are benefits for human readers, like facilitating learning. However, prior research suggests that extensive diacritization can inadvertently impede skilled reading by increasing the required processing time \@BBOPcite\@BAP\@BBNTaha (2016); Ibrahim (2013); Abu-Leil et al. (2014); Midhwah and Alhawary (2020); Roman and Pavard (1987); Hermena et al. (2015)\@BBCP. Nevertheless, diacritics are important morphological markers, even when excessive, and may benefit automated systems in language modeling, machine translation (MT), part-of-speech tagging, morphological analysis, acoustic modeling for speech recognition, and text-to-speech synthesis. For an example in MT, see \@BBOPcitet\@BAP\@BBNFadel et al. (2019); habash-etal-2016-exploiting; diab-etal-2007-arabic\@BBCP.
@@ 1.1 Contributions
We propose a novel method for Partial Diacritization (PD) which seamlessly utilizes existing Arabic FD systems. We exploit a statistical property of Arabic words wherein readers can guess the correct reading of most unmarked words with minimal context. To select the letters to mark, CCPD sees each word twice: (1) within its sentence context, and (2) as an isolated input with no context. The two predictions are combined to retain only those diacritics which present comparatively new information that may aid reading comprehension (Section 4).
@@ 4.2 What Context Do Models See?
Where contains all the words in a segment as used in training D2/TD2.
Contextual Prediction.
Let use and for word and letter contexts as follows:
Where refers to the context around .
@@
sees only the current word at position . Notice the instead of .
We use “word” to refer to syntactically separated
@@ 4.3 Context-Contrastive Partial Diacritization
Using and probabilities, for a word and character , the system outputs a diacritic according to the following functions:
In mode, the characters where applications with and without context agree are not assigned diacritics (regardless of the correctness of the prediction; unknown during inference). Otherwise, the prediction of the contextual model is returned.
In mode, only the characters where the logit of ’s prediction exceeds the logit of ’s by a margin receive a diacritical mark.
@@ 5.1 SR: Selection Rate
The proportion of characters assigned a diacritic by Function (5). Literature shows partial diacritization hovers around – coverage in some professionally published books \@BBOPcitep\@BAP\@BBN(Hallberg, 2022)\@BBCP, while other research suggests that deliberate partial diacritization by native speakers results in a rate around – \@BBOPcitep\@BAP\@BBN(Esmail et al., 2022)\@BBCP. SR is:
[…]
Logical expressions like are: if true, and otherwise.
5.2 Scoped Diacritic Error Rates: P-DER, B-DER
DER in traditional literature is the ratio of diacritics erroneously predicted over all eligible letters in a corpus. Trivially, we parameterize DER by the system and corpus (subset) such that:
@@ 5.2.1 P-DER: Partial Diacritic Error Rate
We calculate the DER on : All characters in the corpus which were assigned diacritics by the system. The predictions come from , following Equation 5. P-DER is:
@@ 5.2.2 B-DER: Basic Diacritic Error Rate
We calculate the DER on the whole corpus for the predictions of . By the intuition in Section 4.3 regarding , we use B-DER as a proxy for the expected human error on plain unmarked text.
[…]
5.3 R-DER: Reader Diacretics Error Rate
We combine the Partial DER and Hidden DER into one general measure of the total expected error when a human reads partially diacritized text. The Hidden DER indicator is proxy for the expected human error on undiacritized text, while Partial DER reports the actual model error on the diacritized text subset. We weigh both numbers by SR and report the Reading-DER (R-DER) as:
5.4 Redri: Reading-DER Improvement
To quantify the expected improvement on reading quality when adding the selected subset of diacritics by some system , we report the ratio of R-DER (proxy for general reader error with PD) to the DER using non-contextual models (B-DER). Thus, Reading-DER Improvement (Redri) is:
System | SR | P-DER | R-DER | Redri | DER | WER | |
\@BBOPcitep\@BAP\@BBNBarqawi, 2017\@BBCP | 8.6% | 17.9% | 3.6%** | 63.2% | 3.58% | 11.19% | |
\@BBOPcitep\@BAP\@BBNFadel et al., 2019\@BBCP (big)* | 8.9% | 7.8% | 1.6%** | 82.8% | 1.60% | 5.08% | |
\@BBOPcitep\@BAP\@BBNAlKhamissi et al., 2020\@BBCP - D2 | 6.5% | 11.2% | 1.8%** | 74.8% | 1.85% | 5.53% | |
—Ours— | (SP, hard) | 6.8% | 15.0% | 2.0%** | 71.8% | 2.00% | 6.42% |
(SP, soft ) | 5.9% | 3.1% | 4.3% | 39.4% | |||
(SP, soft ) | 11.5% | 4.7% | 2.8% | 60.6% | |||
(SP, soft ) | 15.2% | 5.1% | 2.3% | 67.6% | |||
(SP, soft ) | 21.8% | 4.9% | 2.0% | 71.8% | |||
TD2 (MV, hard) | 24.6% | 5.5% | 2.4%** | 90.3% | 2.44% | 7.68% |
@@ 6.1 TD2: Transformer D2
The TD2 model is a Transformer adaptation of D2 \@BBOPcitep\@BAP\@BBN(AlKhamissi et al., 2020)\@BBCP. It comprises two encoders of 2 layers each, for tokens and characters. Both models use the same architecture. The feature and intermediate widths are (768, 2304). The token model is initialized from the pretrained weights of layers 1-2 of CAMeL-Lab/bert-base-arabic-camelbert-mix-ner \@BBOPcitep\@BAP\@BBN(Inoue et al., 2021)\@BBCP as provided by HuggingFace222HuggingFace: \@BBOPcitet\@BAP\@BBNWolf et al. (2020)\@BBCP, while the character model is initialized from layers 3-4.
6.2 Single-Pass Inference
For recurrent (non-Transformer) models like D2, we also test inference without segmentation and majority voting wherein each sentence is passed once and in full to the model (SP). This can save inference time (by avoiding overlapping windows), but sacrifices some reliability (as the model is trained on smaller window sizes).
@@ 7.1 Full Diacritization Performance
Following prior work, we report the Diacritic Error Rate (DER) and Word Error Rate (WER) twice in Table 3, once including case-endings (CE) and once excluding them. CE denotes the diacritics located at the word’s end, usually determined by the word’s syntactic role. Predicting these diacritics is more challenging compared to core-word diacritics (CW), which specify lexical choices and are found elsewhere within the word.
@@ 7.2 Partial Diacritization Performance
Using our proposed indicators: Selection Rate, Partial-DER, Reading-DER, and Reading Improvement, we can highlight some properties of the test models and their interactions with the test sets.
On Tashkeela, we notice that the Recurrent D2 is conservative in its selection with SR at 6.46% versus TD2’s 24.61%. Despite the larger portion of non-diacritized text, its non-contextual error is lower with at 1.14% (versus 1.42%). The low unmarked letter DER numbers suggest that most of the characters skipped by both systems should be guessable for a human reader. Notably, TD2 had many contrastive disagreements leading to its high SR. The lower P-DER from D2 and TD2 (as low as 3.1%–5.5%) at relatively high SR (around 15%; within the rates reported in Section 5.1).
TD2 scores the highest Redri at 90.3%. Thus out of all errors made by over the entire corpus, 90.3% were fixed by TD2 when it output diacritics covering only 24.6% of the corpus. Assuming that the selected letters are the hardest to guess for a reader, this suggests that every tested model is capable of disambiguating the text by 39% to 90% by marking only 5.9% to 24.6% of the letters.
@@ 8 Behavioral Experiment on Partial Diacritization
To test this hypothesis, we conduct a behavioral experiment using machine-predicted partial diacritization masks. The diacritical marks themselves are the true labels. This is to ensure that the presented text is accurate, even if the model output is inaccurate. The question is then: Does CCPD between and work well for choosing the letters to be diacritized? See the results in Figure 4.
@@ 8.1 Behavioral Experiment Setup
We utilize the PsyToolkit online platform333PsyToolkit: \@BBOPcitet\@BAP\@BBNStoet (2010, 2017)\@BBCP to conduct a behavioral experiment aimed at assessing the impact of different levels of diacritization (modes) on reading speed and accuracy. Data was gathered from a group of 15 participants, covering various age groups and native speakers of different Arabic dialects. The majority of participants (10) fell within the 25–40 age group; 2 were below 25, and 3 were above 40. All participants spoke at least 2 languages, while 6 spoke 3 or more. Eight participants spoke either Gulf (4) or Maghrebi (4) Arabic natively, and 7 spoke Egyptian (2), Levantine (2), or Sudanese (2). One did not report a native dialect. All participants reported being native speakers of Arabic.
@@
To assess the impact of diacritics, we select 30 sentences from various domains in the WikiNews testset \@BBOPcitep\@BAP\@BBN(Darwish et al., 2017)\@BBCP. Each sentence is presented in three variants with Zero/Partial/Full diacritics. The Partial variant is produced using our CCPD algorithm and the D2-SP model in selection mode to mask out easy or guessable ground truth diacritics.
@@
The data is split into 3 buckets of 10 sentences each. Each bucket has 3 versions as outlined previously. We create 3 test sets such that the buckets are rotated and each appears exactly once in any mode (Zero/Partial/Full). For example, the first bucket would appear in the test sets A, B, C in its Full, Zero, and Partial variants. This is done to ensure that any participant sees a sentence exactly once, to avoid biasing their rating by repeated exposure. This ensures also that each sentence is seen equal times in each mode.
@@
Participants are shown each sentence for a few seconds proportional to the word count of the sentence: . Then the sentence is hidden and participants are prompted to rate their comprehension by a score between 1 and 5, where 1 indicates difficulty and 5 indicates ease in understanding the sentence. The design intention is to measure the reading experience of a native speaker when scanning a non-technical text in Modern Standard Arabic.
@@ 8.2 Findings
Figure 4 illustrates the aggregated self-reported reading comprehension scores, normalized by the score of the no diacritization (Zero) mode, for each diacritization mode. The results indicate that Full diacritization hinders reading accuracy when participants are provided with a limited amount of time to read the text, aligning with findings discussed in Section 2. Moreover, the Partial diacritization mode sometimes enhances reading comprehension performance compared to the Zero mode, as it can help resolve ambiguities in the text.
The data from this behavioral experiment will be available on the paper’s GitHub repository444github.com/munael/ccpd-partial-arabic-diacritization.
@@ 9 Related Work
In this section, we present the extant body of literature split into two primary categories. The first category centers on the Full Diacritization of Arabic text, encompassing the task of assigning a diacritic label to each individual character within the text. The second category pertains to Partial Diacritization, a distinct approach that seeks to augment reading comprehension by incorporating only the minimal requisite diacritics. The majority of the existing literature belong to the former, while few works focused on the latter \@BBOPcite\@BAP\@BBNAlmanea (2021); Mijlad and El Younoussi (2022)\@BBCP. Other works take a morphological analysis approach \@BBOPcitep\@BAP\@BBN(habash-rambow-2007-arabic; obeid-etal-2020-camel; obeid-etal-2022-camelira; shahrour-etal-2015-improving)\@BBCP.
@@ 9.1 Full Diacritization
Recent literature has focused more on neural-based systems. \@BBOPcitet\@BAP\@BBNBelinkov and Glass (2015)\@BBCP were the first to show that recurrent neural models are suitable candidates to learn the task entirely from data without resorting to manually engineered features such as morphological analyzers and POS taggers. \@BBOPcitet\@BAP\@BBNAlKhamissi et al. (2020)\@BBCP used a hierarchical, BiLSTM-based model that operates on words and characters separately with a cross-level attention connecting the two—enabling SOTA task performance and faster training and inference compared to traditional models. Several other papers as well used recurrent-based models for the automatically diacritizating Arabic text \@BBOPcite\@BAP\@BBNGheith Abandah (2020); Fadel et al. (2019); Boudchiche et al. (2017); Moumen et al. (2018); Al-Thubaity et al. (2020); Darwish et al. (2020)\@BBCP, others used the Transformer architecture \@BBOPcite\@BAP\@BBNMubarak et al. (2019)\@BBCP.

### v3

# A Context-Contrastive Inference Approach To Partial Diacritization

@@ Abstract
Diacritization plays a pivotal role in improving readability and disambiguating the meaning of Arabic texts. Efforts have so far focused on marking every eligible character (Full Diacritization). Comparatively overlooked, Partial Diacritzation (PD) is the selection of a subset of characters to be annotated to aid comprehension where needed. Research has indicated that excessive diacritic marks can hinder skilled readers—reducing reading speed and accuracy. We conduct a behavioral experiment and show that partially marked text is often easier to read than fully marked text, and sometimes easier than plain text. In this light, we introduce Context-Contrastive Partial Diacritization ()—a novel approach to PD which integrates seamlessly with existing Arabic diacritization systems. processes each word twice, once with context and once without, and diacritizes only the characters with disparities between the two inferences. Further, we introduce novel indicators for measuring partial diacritization quality to help establish this as a machine learning task. Lastly, we introduce TD2, a Transformer-variant of an established model which offers a markedly different performance profile on our proposed indicators compared to all other known systems.111Demo: https://huggingface.co/spaces/bkhmsi/Partial-Arabic-Diacritization
A Context-Contrastive Inference
Approach To Partial Diacritization
@@ 1 Introduction
There are benefits for human readers, like facilitating learning. However, prior research suggests that extensive diacritization can inadvertently impede skilled reading by increasing the required processing time \@BBOPcite\@BAP\@BBNTaha (2016); Ibrahim (2013); Abu-Leil et al. (2014); Midhwah and Alhawary (2020); Roman and Pavard (1987); Hermena et al. (2015)\@BBCP. Nevertheless, diacritics are important morphological markers, even when excessive, and may benefit automated systems in language modeling, machine translation (MT), part-of-speech tagging, morphological analysis, acoustic modeling for speech recognition, and text-to-speech synthesis. For an example in MT, see \@BBOPcitet\@BAP\@BBNFadel et al. (2019); Habash et al. (2016); Alqahtani et al. (2016); Diab et al. (2007)\@BBCP.
@@ 1.1 Contributions
We propose a novel method for Partial Diacritization (PD) which seamlessly utilizes existing Arabic FD systems. We exploit a statistical property of Arabic words wherein readers can guess the correct reading of most unmarked words with minimal context. To select the letters to mark, sees each word twice: (1) within its sentence context, and (2) as an isolated input with no context. The two predictions are combined to retain only those diacritics which present comparatively new information that may aid reading comprehension (Section 4).
@@ 4.2 What Context Do Models See?
Let the sub-sequence contain all the words in a segment of length s.t. as used in training D2/TD2.
@@
Thus refers to a segment containing the word , while refers to the intra-word context for letter in word .
Contextual Prediction
sees the full segment.
@@
sees only the current word222Sentences are split on spaces. Words are not tokenized. at position . Note the instead of .
Thus the two modes are defined as follows: is single-word application, while uses sentence context. Both use the same model .
@@ 4.3 Context-Contrastive Partial Diacritization
Using and predictions for a word and letter , the system assigns or omits a diacritic according to :
In hard or disagreement mode, the letters where inferences with and without context agree are left bare (regardless of the correctness of the prediction; unknown during inference). Otherwise, the contextual prediction is returned.
In soft or confidence mode, only the letters where the logit exceeds the logit by a margin receive a diacritical mark.
@@ 5 Performance Indicators
Indicators may use the marked () and unmarked () subsets of the corpus ().
selected | (8) | ||||
unmarked |
@@ 5.1 SR: Selection Rate
SR is the proportion of characters assigned a diacritic by Function (LABEL:eq:pdd-assign). Literature shows partial diacritization hovers around – coverage in some professionally published books \@BBOPcitep\@BAP\@BBN(Hallberg, 2022)\@BBCP, while other research suggests that deliberate partial diacritization by native speakers results in a rate around – \@BBOPcitep\@BAP\@BBN(Esmail et al., 2022)\@BBCP.
5.2 Scoped Diacritic Error Rates DER(f, D)
@@ 5.2.2 B-DER: Basic Diacritic Error Rate
By the intuition in Section 4.3 regarding the non-contextual predictions of , we calculate the DER on the whole corpus and use B-DER as a proxy for human error on plain unmarked text.
5.3 RE-DER: Reader DER
We combine the Partial DER and Basic DER into one general measure of the total error in a partially diacritized text. The Basic DER indicator is used as a proxy which may be taken as an upper bound for mistakes made by an experienced reader. Formally, this indicator incorporates the error of the non-contextual model for the unmarked text, and of the contextual-model for the marked text (because the system explicitly annotates it). Thus RE-DER (Reader DER) is:
5.4 SU: Signal Utilization
How well does utilize its marked subset to disambiguate the text? We gauge how informative the added diacritics are by measuring the change in DER within the marked subset of text. SUspans where positive values denote improvement over no annotation. SU is defined as:
E.g. an SU=50% shows that half of the produced annotations clarify ambiguities (lower RE-DER).
System | SR | SU | P-DER | RE-DER | B-DER | DER | WER | |
Reference Range | ||||||||
\@BBOPcitep\@BAP\@BBNBarqawi, 2017\@BBCP | 8.6 | 70.9 | 17.9 | †3.6 | 9.7 | 3.58 | 11.19 | |
\@BBOPcitep\@BAP\@BBNFadel et al., 2019\@BBCP (big)* | 8.9 | 86.5 | 7.8 | †1.6 | 9.3 | 1.60 | 5.08 | |
\@BBOPcitep\@BAP\@BBNAlKhamissi et al., 2020\@BBCP - D2 | 6.5 | 81.5 | 11.2 | †1.8 | 7.1 | 1.85 | 5.53 | |
—Ours— | D2 (SP, hard) | 6.8 | 75.0 | 15.0 | †2.0 | 7.1 | 2.00 | 6.42 |
D2 (SP, soft ) | 5.9 | 47.5 | 3.1 | 4.3 | ||||
D2 (SP, soft ) | 11.5 | 37.4 | 4.7 | 2.8 | ||||
D2 (SP, soft ) | 15.2 | 31.6 | 5.1 | 2.3 | ||||
D2 (SP, soft ) | 21.8 | 23.4 | 4.9 | 2.0 | ||||
TD2 (MV, hard) | 24.6 | 91.5 | 5.5 | †2.4 | 25.0 | 2.44 | 7.68 |
@@ 6.1 TD2: Transformer D2
The TD2 model is a Transformer adaptation of D2 \@BBOPcitep\@BAP\@BBN(AlKhamissi et al., 2020)\@BBCP. It comprises two encoders of 2 layers each, for tokens and characters. Both models use the same architecture. The feature and intermediate widths are (768, 2304). The token model is initialized from the pretrained weights of layers 1-2 of CAMeL-Lab/bert-base-arabic-camelbert-mix-ner \@BBOPcitep\@BAP\@BBN(Inoue et al., 2021)\@BBCP as provided by HuggingFace333HuggingFace: \@BBOPcitet\@BAP\@BBNWolf et al. (2020)\@BBCP. The character model is initialized from layers 3-4 of the same pretrained weights.
6.2 SP: Single-Pass Inference
For recurrent (non-Transformer) models like D2, we also test inference without segmentation and majority voting. Each sentence is passed in full the model, once. This can save inference time (by avoiding overlapping windows), but sacrifices some reliability (as the model had been trained on smaller window sizes).
@@ 7.1 Full Diacritization Performance
Following prior work, we report the Diacritic Error Rate (DER) and Word Error Rate (WER) including case-endings which are located at the word’s end, usually determined by the word’s syntactic role. Predicting these diacritics is more challenging compared to core-word diacritics, which specify lexical properties and lie elsewhere within the word.
@@ 7.2 Partial Diacritization Performance
Intuition
Overall it is desirable to observe: (1) a natural SR value close to native human annotation rates; (2) a low B-DER close to the SR, signifying a capable model which allows a clean partial annotation of only the hardest letters, (3) a low P-DER to signify high accuracy in the committed diacritics, and (4) a low RE-DER which gauges the overall expected reading experience given the partial annotation and the expected guessing error. Overall, this corresponds to a balance between a high SU, a natural SR, and a low P-DER, roughly ordered.
Model results
Notice that the recurrent D2 is conservative in its selection with SR at 6.46% versus TD2’s 24.61%.
TD2 under-performs in non-contextual mode with B-DER at 25% compared to 7.1% by D2. Counter-intuitively, this leads it to a higher SU at 91.5% as its contextual mode corrects 22.5 of every 25 errors its non-contextual mode commits.
Both systems agree that much of the text (at least 75%) is reasonably guessable by the reader without annotation, by taking B-DER as an upper-bound on human guess error.
We may claim that between 75% and 93% of characters require no diacritization to disambiguate (the extremes of the two systems).
Both D2 and TD2 correct most of their B-DER error via , implying that most letters which had errors in non-contextual mode were accurately selected by and correctly predicted by , leading to an RE-DER < B-DER.
Soft marking
SR can be tweaked in soft selection mode. Let’s analyze the model with SR=15.2% (Table 3) which provides a low P-DER at 5.1% (few diacritics added are erroneous) and a low RE-DER at 2.3% (around 4.8 out of 7.1% expected errors are corrected via the added diacritics). Notice the low SU (utilization) however, as only 31.6% of the annotated letters contribute to the drop in error.
Assuming that the selected letters are the hardest to guess for a reader, this suggests that D2-SP models can disambiguate the text by 39.4% to 71.8% (relative improvement in RE-DER/B-DER) by marking only 5.9% to 21.8% of it.
While RE-DER improves, P-DER and SU worsen, suggesting that increased annotation may give only marginal improvement in overall readability if the final model output is not perfectly clean.
@@ 8 Behavioral Experiment on Partial Diacritization
To test this hypothesis, we conduct a behavioral experiment using machine-predicted partial diacritization masks. The diacritical marks themselves are the true labels. This is to ensure that the presented text is accurate, even if the model output is sub-optimal. The question is then: Does using and select useful letters to be diacritized? See the results in Figure 4.
@@ 8.1 Behavioral Experiment Setup
We utilize the PsyToolkit online platform444PsyToolkit: \@BBOPcitet\@BAP\@BBNStoet (2010, 2017)\@BBCP to conduct a behavioral experiment aimed at assessing the impact of different levels of diacritization (modes) on reading speed and accuracy. Data was gathered from a group of 15 participants, covering various age groups and native speakers of different Arabic dialects. The majority of participants (10) fell within the 25–40 age group; 2 were below 25, and 3 were above 40. All participants spoke at least 2 languages, while 6 spoke 3 or more. Eight participants spoke either Gulf (4) or Maghrebi (4) Arabic natively, and 7 spoke Egyptian (2), Levantine (2), or Sudanese (2). One did not report a native dialect. All participants reported being native speakers of Arabic.
@@
To measure the impact of diacritics, we select 30 sentences from various domains in the WikiNews testset \@BBOPcitep\@BAP\@BBN(Darwish et al., 2017)\@BBCP. Each sentence is presented in three variants with Zero/Partial/Full diacritics. The Partial variant is produced using our algorithm and the D2-SP model in selection mode to mask out easy or guessable ground truth diacritics.
@@
The data is split into 3 buckets of 10 sentences each. We create 3 test sets such that the buckets are rotated and each appears exactly once in any mode (Zero/Partial/Full). For example, the first bucket would appear in the test sets A, B, C in its Full, Zero, and Partial variants. This is done to ensure that any participant sees a sentence exactly once, to avoid biasing their rating by repeated exposure. This ensures also that each sentence is seen equal times in each mode.
@@
Participants are shown each sentence for a few seconds proportional to the word count of the sentence: . Then the participants are prompted to rate their comprehension by a score from 1 to 5, where 1 indicates difficulty and 5 indicates ease in understanding the sentence. The design intention is to measure the reading experience of a native speaker when scanning a non-technical text in Modern Standard Arabic.
@@ 8.2 Findings
Figure 4 illustrates the per-mode average self-reported reading comprehension scores, normalized by the score of the no diacritization (Zero) mode. The results indicate that Full diacritization hinders reading accuracy when participants are provided with a limited amount of time to read the text, aligning with findings discussed in Section 2. In contrast, the Partial diacritization mode sometimes enhances reading comprehension performance compared to the Zero mode.
The data from this behavioral experiment will be available on the paper’s GituHub repository555GitHub: munael/arabic-partial-diacritization.
@@ 9 Related Work
The literature has dealth with both types of diacritization: Full (FD) and Partial (PD). The majority has focused on FD. In PD, a distinct approach that seeks to augment reading comprehension by incorporating only the minimal requisite diacritics, we note a few works focus \@BBOPcitep\@BAP\@BBN(Almanea, 2021; Mijlad and El Younoussi, 2022)\@BBCP. Other works take a morphological analysis approach \@BBOPcitep\@BAP\@BBN(Obeid et al., 2022, 2020; Alqahtani et al., 2016; Shahrour et al., 2015; Habash and Rambow, 2007)\@BBCP.
@@ 9.1 Full Diacritization
Recent literature has focused more on neural-based systems. \@BBOPcitet\@BAP\@BBNBelinkov and Glass (2015)\@BBCP were the first to show that recurrent neural models are suitable candidates to learn the task entirely from data without resorting to manually engineered features such as morphological analyzers and POS taggers. \@BBOPcitet\@BAP\@BBNAlKhamissi et al. (2020)\@BBCP used a hierarchical, BiLSTM-based model that operates on words and characters separately with a cross-level attention connecting the two—enabling SOTA task performance and faster training and inference compared to traditional models. Many prior works have used recurrent-based models for the FD task \@BBOPcite\@BAP\@BBNAl-Thubaity et al. (2020); Darwish et al. (2020); Gheith Abandah (2020); Fadel et al. (2019); Boudchiche et al. (2017); Moumen et al. (2018)\@BBCP, others used the Transformer architecture \@BBOPcite\@BAP\@BBNMubarak et al. (2019)\@BBCP, while others used ConvNet architectures \@BBOPcite\@BAP\@BBNAlqahtani et al. (2019b)\@BBCP.